OpenStudy (anonymous):

Which of the following is a candidate for being an extreme value for the function f(x)= -0.5x^2 2^x

OpenStudy (inkyvoyd):

Well, that didn't come out right...

OpenStudy (anonymous):

it is given options of 0, 1/2 , 1 and 2

OpenStudy (anonymous):

but my answer is so off...

OpenStudy (blockcolder):

Is this f(x)? \[f(x)=-0.5x^2 2^x\]

OpenStudy (blockcolder):

If yes, then let's look at it closely. \[x^2 \geq 0\ \text{and}\ 2^x>0\] Therefore, \[x^2 2^x \geq 0 \\ \Rightarrow -0.5x^2 2^x \leq 0\] Now, can any of 1/2, 1, and 2 be a maximum value?

OpenStudy (anonymous):

no if it must be less or equal to 0 ! :)

OpenStudy (anonymous):

but dont u need to find the first derivative ?

OpenStudy (blockcolder):

There's no need, really. It wasn't really asked WHAT the max value was, only which one could possibly be a max value. :D

OpenStudy (anonymous):

oh i get it thanks so much for ur help!

OpenStudy (blockcolder):

No problem. :D